OpenStudy (anonymous):

1: 5/7 of 7/10 2: 2% of 700 3: 1% of 600 4: 6 hours and 54 minutes plus 6 hours and 18 minute Ribbon for even 1 answer!

OpenStudy (anonymous):

1.5/7*7/10 =1/2 2.700/100*2 =14 3.600/100*1 =6 4.13h 12min

OpenStudy (anonymous):

1.1/2 2.14 3.6 4. 13 hours 12 minutes

OpenStudy (anonymous):

Number 4 written as decimal?